# Longitudinal efficient adjustment sets for time-varying treatment effect estimation in nonparametric causal graphical models

### v1

@@ 3 Proposal
In practice, a sufficient time-dependent adjustment set according to Definition 2 can be constructed by first identifying a time-dependent adjustment set according to Definition 1 and exploiting conditional independencies that can be read from the DAG to exclude unnecessary variables from the identification formula. Algorithm 1 describes a process for achieving this.
@@
1. Identify a sufficient time-dependent adjustment set $\mathbf{Z}$ according to Definition 1 using the SWIG associated with the initial DAG. Express this sufficient time-dependent adjustment set using the notation of Definition 2. In the following steps, the DAG (and not the SWIG) will be used will be used to read conditional independencies.
2. Let $Q_{p+1}=Y$.
for $k=p$ to 1 do
     3.1 Read from the DAG conditional independencies between $Q_{k+1}$ and variables in $\{A_{k},\mathbf{Z}_{k}\}$ and exclude variables from $\mathbf{Z}_{k}$ that are conditionally independent. Denote as $\mathbf{G}_{k}$ the set of variables that are retained.
     3.2 Construct a modified DAG with an additional $Q_{k}$ node and remove node $Q_{k+1}$. Add arrows pointing into $Q_{k}$ stemming from the variables $\mathbf{G}_{k}$ that were retained. Remove the node $A_{k}$ from the modified SWIG.
@@
We now return to Example 1 to illustrate this process.
@@
Step 1 of the algorithm was achieved in the first part of Example 1, where the sufficient time-dependent adjustment set $\mathbf{Z}=(\mathbf{Z_{0}}=C_{0,2},\mathbf{Z_{1}}=C_{1,2})$ was found. Using the notation introduced in Definition 2, this sufficient time-dependent adjustment set can be expressed as ${\mathbf{Z}=(\mathbf{Z}_{0}=C_{0,2},\mathbf{Z}_{1}=\{C_{1,2},A_{0},C_{0,2}\})}$, which leads to the same identification formula as previously, that is
$\mathbb{E}_{P}(Y^{\mathbf{a}})=\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid A_{1}=a_{1},C_{1,2},A_{0}=a_{0},C_{0,2})\mid A_{0}=a_{0},C_{0,2}\}].$
For $k=p=1$, in Step 2, we let $Q_{2}=Y$. The resulting DAG is the same as the one in Figure 1 (a), with $Y$ replaced by $Q_{2}$. In Step 3.1, by inspecting this DAG, it can be seen that $Q_{2}\perp\!\!\!\perp C_{0,2}\mid A_{1},C_{1,2},A_{0}$ or equivalently that $Y\perp\!\!\!\perp C_{0,2}\mid A_{1},C_{1,2},A_{0}$. As such,
$\mathbb{E}_{p}(Y^{\mathbf{a}})$ $=\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid A_{1}=a_{1},C_{1,2},A_{0}=a_{0},C_{0,2})\mid A_{0}=a_{0},C_{0,2}\}]$
$=\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid A_{1}=a_{1},C_{1,2},A_{0}=a_{0})\mid A_{0}=a_{0},C_{0,2}\}]$
and we deduce that $(\mathbf{Z}_{0}=C_{0,2},\mathbf{G}_{1}=\{C_{1,2},A_{0}\})$ is a sufficient time-dependent adjustment set according to Definition 2. In Step 3.2, we produce the modified DAG in Figure 2 with $Q_{1}=\mathbb{E}_{P}(Y\mid A_{1}=a_{1},C_{1,2},A_{0}=a_{0})$ and deleted $Y$.
[…]
Figure 2: The modified causal directed acyclic graph (DAG) obtained in Step 3.2 of Example 1
For $k=0$, returning to Step 3.1, we can see from Figure 2 that $C_{0,2}$ is associated with $Q_{1}$ conditional on $A_{0}$. No additional simplification is thus possible. The resulting time-dependent adjustment set according to Definition 2 is $\mathbf{G}=(\mathbf{G}_{0}=C_{0,2},\mathbf{G}_{1}=\{C_{1,2},A_{0}\})$. Remark that this time-dependent adjustment set is not admissible according to Definition 1.
In the next section, we turn our attention to comparing the asymptotic variance of nonparametric efficient estimators of causal contrasts $\Delta(P;\mathcal{G})\equiv\sum_{\mathbf{a}\in\mathcal{A}}c_{\mathbf{a}}\mathbb{E}_{P}(Y^{\mathbf{a}})$ based on different sufficient time-dependent adjustment sets according to our proposed definition.
4 Comparison of time-dependent adjustment sets
@@
$b_{a_{k}}(\mathbf{Z}_{k};P)\equiv\mathbb{E}_{P}[...\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid\mathbf{a}_{p},\mathbf{Z}_{p}(\bar{a}_{p-1}))\mid\mathbf{a}_{p-1},\mathbf{Z}_{p-1}(\bar{a}_{p-2})\}...\mid\mathbf{a}_{k},\mathbf{Z}_{k}(\bar{a}_{k-1})],$
$\pi_{a_{k}}(\mathbf{Z}_{k};P)\equiv P(A_{k}=a_{k}\mid\mathbf{Z}_{k}(\bar{a}_{k-1})),$
$\psi_{P,\mathbf{a}}(\mathbf{Z};\mathcal{G})=\dfrac{I_{\bar{\mathbf{a}}_{p}}(\bar{\mathbf{A}}_{p})}{\lambda_{\bar{\mathbf{a}}_{p}}(\bar{\mathbf{Z}}_{p};P)}\{Y-b_{\mathbf{a}_{p}}(\mathbf{Z}_{p};P)\}+\displaystyle{\sum_{k=0}^{p}}\dfrac{I_{\bar{\mathbf{a}}_{k-1}}(\bar{\mathbf{A}}_{k-1})\{b_{\mathbf{a}_{k}}(\mathbf{Z}_{k};P)-b_{\mathbf{a}_{k-1}}(\mathbf{Z}_{k-1};P)\}}{\lambda_{\bar{\mathbf{a}}_{k-1}}(\bar{\mathbf{Z}}_{k-1};P)},$
@@
where $b_{a_{-1}}(\mathbf{Z}_{-1};P)\equiv\chi_{\mathbf{a}}(P;\mathcal{G})$, $\lambda_{\bar{\mathbf{a}}_{k-1}}(\bar{\mathbf{Z}}_{k-1};P)\equiv\displaystyle{\prod_{j=0}^{k-1}}\pi_{a_{j}}(\mathbf{Z}_{j};P)$ and $I_{\bar{\mathbf{a}}_{-1}}(\bar{\mathbf{A}}_{-1})\{\lambda_{\bar{\mathbf{a}}_{-1}}(\bar{\mathbf{Z}}_{-1};P)\}^{-1}\equiv 1$. The quantity $\psi_{P,\mathbf{a}}(\mathbf{Z};\mathcal{G})$ is the efficient influence function of the counterfactual expectation $\mathbb{E}_{P}(Y^{\mathbf{a}})$ under the nonparametric model using the time-dependent adjustment set $\mathbf{Z}$ [9]. Similar to the Cramér-Rao bound, the variance of this efficient influence function scaled by a factor $1/n$ defines a lower bound for the asymptotic variance of nonparametric estimators of $\mathbb{E}_{P}(Y^{\mathbf{a}})$ for a given time-dependent adjustment set [10, 11]. We further define $\sigma^{2}_{\mathbf{a},\mathbf{Z}}(P)$ and $\sigma^{2}_{\Delta,\mathbf{Z}}(P)$ as the asymptotic variance of a nonparametric efficient estimator of the counterfactual mean $\chi_{\mathbf{a}}(P;\mathcal{G})$ and the causal contrast $\Delta(P;\mathcal{G})$ under the time-dependent adjustment set $\mathbf{Z}$. Note that the latter can be obtained from the former using the functional delta method [12] or tricks delineated in [13]. Using this notation, the following two lemmas compare the asymptotic variance of nonparametric efficient estimators of causal contrasts between nested sufficient time-dependent adjustment sets.
Lemma 4.1 (Inclusion of additional variables $\mathbf{G}$ in a sufficient time-dependent adjustment set $\mathbf{Z}=\mathbf{B}$).
Let $\mathbf{K}=(\mathbf{K}_{0},\mathbf{K}_{1},...,\mathbf{K}_{p})$ be a sufficient time-dependent adjustment set according to Definition 1 and let ${\mathbf{B}_{k}\subseteq\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}}$ and ${\mathbf{G}_{k}\subseteq\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}\setminus\mathbf{B}_{k}},k=0,...,p$ be two disjoint subsets of $\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}$ such that $(\mathbf{G}_{k},\mathbf{B}_{k})=\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}$. Suppose $\mathbf{B}=(\mathbf{B}_{0},\mathbf{B}_{1},...,\mathbf{B}_{p})$ is a sufficient time-dependent adjustment set according to Definition 2 that satisfies
$A_{k}\perp\!\!\!\perp\mathbf{G}_{k}\ \mid\ \mathbf{B}_{k},\ for\ k=0,...,p.$ (3)
@@
Then $(\mathbf{G},\mathbf{B})=\{(\mathbf{G}_{0},\mathbf{B}_{0}),(\mathbf{G}_{1},\mathbf{B}_{1}),...,(\mathbf{G}_{p},\mathbf{B}_{p})\}$ is also a sufficient time-dependent adjustment set according to Definition 2 and
$\sigma^{2}_{\mathbf{a},\mathbf{B}}(P)-\sigma^{2}_{\mathbf{a},\mathbf{G},\mathbf{B}}(P)\geq 0\ \ and\ \ \sigma^{2}_{\Delta,\mathbf{B}}(P)-\sigma^{2}_{\Delta,\mathbf{G},\mathbf{B}}(P)\geq 0.$
Lemma 4.2 (Exclusion of variables $\mathbf{B}$ from a sufficient time-dependent adjustment set $\mathbf{Z}=(\mathbf{G},\mathbf{B})$).
Let $\mathbf{K}=(\mathbf{K}_{0},\mathbf{K}_{1},...,\mathbf{K}_{p})$ be a sufficient time-dependent adjustment set according to Definition 1 and let ${\mathbf{B}_{k}\subseteq\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}}$ and ${\mathbf{G}_{k}\subseteq\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}\setminus\mathbf{B}_{k}},k=0,...,p$ be two disjoint subsets of $\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}$ such that $(\mathbf{G}_{k},\mathbf{B}_{k})=\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}$. Suppose $(\mathbf{G},\mathbf{B})=\{(\mathbf{G}_{0},\mathbf{B}_{0}),(\mathbf{G}_{1},\mathbf{B}_{1}),...,(\mathbf{G}_{p},\mathbf{B}_{p})\}$ is a sufficient time-dependent adjustment set according to Definition 2 that satisfies
$Y\perp\!\!\!\perp\mathbf{B}_{p}\ \mid\ \mathbf{G}_{p},\mathbf{A}_{p}$ (4)

### v2

@@ 3.2 Construction of a sufficient time-dependent adjustment set according to Definition 2
In practice, a sufficient time-dependent adjustment set according to Definition 2 can be constructed by first identifying a time-dependent adjustment set according to Definition 1 and exploiting conditional independencies that can be read from the DAG to exclude unnecessary variables from the identification formula. Algorithm 1 below provides a structured procedure for achieving this. This algorithm employs a recursive procedure, where the problem of estimating the counterfactual expectation under a time-dependent treatment regime $\boldsymbol{A}=\boldsymbol{a}$ is represented as a nested sequence of estimation problems where each problem entails estimating a counterfactual expectation under a given point treatment. Before presenting Algorithm 1, we present an example of this process with theoretical arguments supporting each step.
To simplify the presentation, we consider a two time point setting, but a similar recursive procedure can be applied for any number of time points. Let $\boldsymbol{K}=(\boldsymbol{K}_{0},\boldsymbol{K}_{1})$ be a sufficient adjustment set according to Definition 1 and $\mathcal{G}$ be a causal graph for the distribution $P$ of $\{\boldsymbol{X},\boldsymbol{A},Y\}$. The counterfactual mean can thus be expressed as
$\mathbb{E}_{P}\left(Y^{\boldsymbol{a}}\right)$ $=\mathbb{E}_{P}\!\left[\mathbb{E}_{P}\!\left\{\mathbb{E}_{P}\!\left(Y\,\middle|\,\boldsymbol{A}=\boldsymbol{a},\bar{\boldsymbol{K}}_{1}\right)\,\middle|\,A_{0}=a_{0},\boldsymbol{K}_{0}\right\}\right].$ (3)
This expression highlights two nested estimation problems, one for each of the inner conditional expectations.
The first estimation problem is
$\mathbb{E}_{P}\!\left(Y\,\middle|\,\boldsymbol{A}=\boldsymbol{a},\bar{\boldsymbol{K}}_{1}\right).$
Let $\boldsymbol{B}_{1}$ and $\boldsymbol{G}_{1}$ be a partition of $\{A_{0},\boldsymbol{K}_{0},\boldsymbol{K}_{1}\}$ such that $Y\perp\!\!\!\perp_{\mathcal{G}}\boldsymbol{B}_{1}|\,\boldsymbol{G}_{1},A_{1}$. Then this conditional expectation can be written as
$\mathbb{E}_{P}(Y\mid\boldsymbol{A}=\boldsymbol{a},\bar{\boldsymbol{K}}_{1})=\mathbb{E}_{P}(Y\mid A_{1}=a_{1},\boldsymbol{G}_{1}).$
Remark that if $A_{0}$ is part of $\boldsymbol{G}_{1}$, it is fixed at $A_{0}=a_{0}$. As such, the problem of estimating $\mathbb{E}_{P}(Y\mid\boldsymbol{A}=\boldsymbol{a},\bar{\boldsymbol{K}}_{1})$ entails estimating a counterfactual expectation under the point-treatment $A_{1}=a_{1}$ for each possible value that the variables in $\boldsymbol{G}_{1}$ can take.
Defining $Q_{0}(a_{1},\boldsymbol{G}_{1})\equiv\mathbb{E}_{P}(Y\mid A_{1}=a_{1},\boldsymbol{G}_{1})$, which we simply denote as $Q_{0}$ when no ambiguity arises, the second estimation problem can be written as
$\mathbb{E}(Q_{0}\mid A_{0}=a_{0},\boldsymbol{K}_{0}).$
Let $\boldsymbol{B}_{0}$ and $\boldsymbol{G}_{0}$ be disjoint subsets of $\boldsymbol{K}_{0}$ such that $Q_{0}\perp\!\!\!\perp_{\mathcal{G}}\boldsymbol{B}_{0}\mid\boldsymbol{G}_{0}$. The second estimation problem can therefore be written as
$\mathbb{E}_{P}\!\left(Q_{0}\,\middle|\,A_{0}=a_{0},\boldsymbol{G}_{0}\right).$
This again corresponds to the estimation of the counterfactual expectation under a point-treatment.
A key step in the above procedure is the correct identification of $\boldsymbol{B}_{0}$ such that $Q_{0}\perp\!\!\!\perp_{\mathcal{G}}\boldsymbol{B}_{0}\mid\boldsymbol{G}_{0}$. To simplify this task, we can construct a subgraph $\mathcal{G}_{0}$ based on the initial DAG $\mathcal{G}$. The idea is to simplify the structure of the initial DAG to focus on the problem of evaluating $Q_{0}\perp\!\!\!\perp_{\mathcal{G}}\boldsymbol{B}_{0}\mid\boldsymbol{G}_{0}$ without losing any information in the initial graph that is relevant to this problem. These subgraphs will also play an important role in identifying optimal time-dependent adjustment sets. The first step of the construction of $\mathcal{G}_{0}$ is to remove the node $Y$ (since it is no longer relevant) and to add a node $Q_{0}$. For a given $a_{1}$, remark that $Q_{0}$ is only a function of $\boldsymbol{G}_{1}$. As such, add arrows from all variables in $\boldsymbol{G}_{1}$ to $Q_{0}$. Next, because $A_{1}$ is fixed at $a_{1}$ for all individuals when calculating $Q_{0}$, no association between other nodes and $Q_{0}$ can pass through $A_{1}$. Consequently, the node $A_{1}$ and all arrows pointing into it or from it are removed from $\mathcal{G}_{0}$. Nodes $\boldsymbol{X}_{1}$ are also superfluous since our interest lies in $\{\boldsymbol{B}_{0},\boldsymbol{G}_{0}\}=\boldsymbol{K}_{0}\subseteq\boldsymbol{X}_{0}$. However, before removing these nodes, we need to verify if some of the association between $Q_{0}$ and $\boldsymbol{X}_{0}$ passes only through $\boldsymbol{X}_{1}$. To do so, we replace any causal path from variables in $\boldsymbol{X}_{0}$ to $Q_{0}$ passing only through variables in $\boldsymbol{X}_{1}$ as a direct arc from these variables to $Q_{0}$. The same process is applied for all causal paths from $A_{0}$ to $Q_{0}$ passing only through variables in $\boldsymbol{X}_{1}$. Variables $\boldsymbol{X}_{1}$ can now be removed from $\mathcal{G}_{0}$. The resulting subgraph can now be used to more easily evaluate relevant d-separation conditions and identify $\boldsymbol{B}_{0}$. We now present Algorithm 1.
@@
1. Identify a sufficient time-dependent adjustment set $\boldsymbol{K}$ according to Definition 1 using the SWIG associated with the initial DAG. Express this sufficient time-dependent adjustment set as $\boldsymbol{Z}$ using the notation of Definition 2. In the following steps, the DAG (and not the SWIG) will be used to read d-separation conditions.
2. Read from the initial DAG d-separation between $Y$ and variables in $\{A_{T},\boldsymbol{Z}_{T}\}$ and exclude variables from $\boldsymbol{Z}_{T}$ that are d-separated from $Y$ by $A_{T}$ and the other variables in $\boldsymbol{Z}_{T}$. Denote as $\boldsymbol{G}_{T}$ the set of variables that are retained.
3. Let $Q_{T}=Y$.
for t=T1t=T-1 to 0 do
  3.1 Construct a subgraph $\mathcal{G}_{t}$ following these steps in order:
  • i)
    remove node $Q_{t+1}$ and add a novel node $Q_{t}$.
  • ii)
    add arrows from all nodes in $\boldsymbol{G}_{t+1}$ to $Q_{t}$.
  • iii)
    remove the node $A_{t+1}$ and all arrows pointing into it or from it.
  • iv)
    Change all causal paths from any variable in $\{\bar{A}_{t},\bar{\boldsymbol{X}}_{t}\}$ to $Q_{t}$ that only pass through variables in $\boldsymbol{X}_{t+1}$ into an arrow from the starting variable to $Q_{t}$, that is, for all $k$, $0\leq k\leq t$, change all paths of the form ${X}_{k\ell}\rightarrow X_{(t+1)j_{1}}\rightarrow\cdots\rightarrow X_{(t+1)j_{m}}\rightarrow Q_{t}$ to ${X}_{k\ell}\rightarrow Q_{t}$, and ${A}_{k}\rightarrow X_{(t+1)j_{1}}\rightarrow\cdots\rightarrow X_{(t+1)j_{m}}\rightarrow Q_{t}$ to ${A}_{k}\rightarrow Q_{t}$ with $\ell\in\{1,\dots,n_{xk}\}$, $j_{1},\dots,j_{m}\in\{1,2,\dots,n_{x(t+1)}\},m\geq 1$.
  • v)
    remove all nodes $\boldsymbol{X}_{t+1}$ and all arrows pointing into them or from them.
  3.2 Read from the subgraph $\mathcal{G}_{t}$ d-separation between $Q_{t}$ and variables in $\{A_{t},\boldsymbol{X}_{t}\}$ and exclude variables from $\boldsymbol{X}_{t}$ that are d-separated from $Q_{t}$ by $A_{t}$ and the other variables in $\boldsymbol{X}_{t}$. Denote as $\boldsymbol{G}_{t}$ the set of variables that are retained.
@@ 3.1 Construct a subgraph $\mathcal{G}_{t}$ following these steps in order:
4. Return $\boldsymbol{G}=(\boldsymbol{G}_{0},\dots,\boldsymbol{G}_{T})$ a sufficient time-dependent adjustment set according to Definition 2.
We return to Example 1 to illustrate the application of Algorithm 1 concretely.
@@
Step 1 of the algorithm was achieved in the first part of Example 1, where the sufficient time-dependent adjustment set $\boldsymbol{K}=(\boldsymbol{K}_{0}=X_{02},\boldsymbol{K}_{1}=X_{12})$ was found. Using the notation introduced in Definition 2, this sufficient time-dependent adjustment set can be expressed as $\boldsymbol{Z}=(\boldsymbol{Z}_{0}=X_{02},\boldsymbol{Z}_{1}=\{X_{12},A_{0},X_{02}\})$}, which leads to the same identification formula as previously, that is
$\mathbb{E}_{P}(Y^{\boldsymbol{a}})=\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid A_{1}=a_{1},X_{12},A_{0}=a_{0},X_{02})\mid A_{0}=a_{0},X_{02}\}].$
In Step 2, by inspecting the DAG, it can be seen that $Y\perp\!\!\!\perp_{\mathcal{G}}X_{02}\mid A_{1},X_{12},A_{0}$. As such,
$\mathbb{E}_{p}(Y^{\boldsymbol{a}})$ $=\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid A_{1}=a_{1},X_{12},A_{0}=a_{0},X_{02})\mid A_{0}=a_{0},X_{02}\}]$
$=\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid A_{1}=a_{1},X_{12},A_{0}=a_{0})\mid A_{0}=a_{0},X_{02}\}]$
and we deduce that $(\boldsymbol{Z}_{0}=X_{02},\boldsymbol{G}_{1}=\{X_{12},A_{0}\})$ is a sufficient time-dependent adjustment set according to Definition 2.
For $t=0$, in Step 3.1, we produce the subgraph $\mathcal{G}_{0}$ in Figure 2. To obtain this graph, the following steps were followed: i) $Y$ has been removed and $Q_{0}=\mathbb{E}_{P}(Y\mid A_{1}=a_{1},X_{12},A_{0}=a_{0})$ has been added, ii) arrows $X_{12}\to Q_{0}$ and $A_{0}\to Q_{0}$ have been added, iii) $A_{1}$ and all arrows pointing into it or from it were removed, iv) an arrow $X_{02}\to Q_{0}$ was added because there was a causal path from it to $Q_{0}$ only passing through variables in $\boldsymbol{X}_{1}$ ($X_{02}\to X_{11}\to X_{12}\to Q_{0}$), and v) the nodes $X_{11}$ and $X_{12}$ were removed.
$X_{01}$$X_{02}$$A_{0}$$Q_{0}$
Figure 2: The subgraph $\mathcal{G}_{0}$ obtained in Step 3.1 of Example 1
In Step 3.2, we can see from Figure 2 that $X_{02}$ is associated with $Q_{0}$ conditional on $A_{0}$. No additional simplification is thus possible. The resulting simplified time-dependent adjustment set according to Definition 2 is $\boldsymbol{G}=(\boldsymbol{G}_{0}=X_{02},\boldsymbol{G}_{1}=\{X_{12},A_{0}\})$. Remark that this time-dependent adjustment set is not admissible according to Definition 1.
In the next section, we turn our attention to comparing the asymptotic variance of nonparametric efficient estimators of causal contrasts $\Delta(P;\mathcal{G})\equiv\sum_{\boldsymbol{a}\in\mathcal{A}}c_{\boldsymbol{a}}\mathbb{E}_{P}(Y^{\boldsymbol{a}})$ based on different sufficient time-dependent adjustment sets according to our proposed definition.
4 Comparison of sufficient time-dependent adjustment sets
@@
$b_{{a}_{t}}(\boldsymbol{Z}_{t};P)\equiv\mathbb{E}_{P}[...\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid{a}_{T},\boldsymbol{Z}_{T}(\bar{\boldsymbol{a}}_{T-1}))\mid{a}_{T-1},\boldsymbol{Z}_{T-1}(\bar{\boldsymbol{a}}_{T-2})\}...\mid{a}_{t},\boldsymbol{Z}_{t}(\bar{\boldsymbol{a}}_{t-1})],$
$\pi_{a_{t}}(\boldsymbol{Z}_{t};P)\equiv P(A_{t}=a_{t}\mid\boldsymbol{Z}_{t}(\bar{\boldsymbol{a}}_{t-1})),$
$\psi_{P,\boldsymbol{a}}(\boldsymbol{Z};\mathcal{G})=\dfrac{I_{\bar{\boldsymbol{a}}_{T}}(\bar{\boldsymbol{A}}_{T})}{\lambda_{\bar{\boldsymbol{a}}_{T}}(\bar{\boldsymbol{Z}}_{T};P)}\{Y-b_{{a}_{T}}(\boldsymbol{Z}_{T};P)\}+\displaystyle{\sum_{t=0}^{T}}\dfrac{I_{\bar{{\boldsymbol{a}}}_{t-1}}(\bar{\boldsymbol{A}}_{t-1})\{b_{{a}_{t}}(\boldsymbol{Z}_{t};P)-b_{{a}_{t-1}}(\boldsymbol{Z}_{t-1};P)\}}{\lambda_{\bar{\boldsymbol{a}}_{t-1}}(\bar{\boldsymbol{Z}}_{t-1};P)},$
where
${b_{a_{-1}}(\boldsymbol{Z}_{-1};P)\equiv\chi_{\boldsymbol{a}}(P;\mathcal{G})},$
$\lambda_{\bar{\boldsymbol{a}}_{t-1}}(\bar{\boldsymbol{Z}}_{t-1};P)\equiv\displaystyle{\prod_{k=0}^{t-1}}\pi_{a_{k}}(\boldsymbol{Z}_{k};P)\text{ and}$
$I_{\bar{\boldsymbol{a}}_{-1}}(\bar{\boldsymbol{A}}_{-1})\{\lambda_{\bar{\boldsymbol{a}}_{-1}}(\bar{\boldsymbol{Z}}_{-1};P)\}^{-1}\equiv 1.$
@@
The quantity $\psi_{P,\boldsymbol{a}}(\boldsymbol{Z};\mathcal{G})$ is the efficient influence function of the counterfactual expectation $\mathbb{E}_{P}(Y^{\boldsymbol{a}})$ under the nonparametric model using the time-dependent adjustment set $\boldsymbol{Z}$ [8]. Similar to the Cramér-Rao bound, the variance of this efficient influence function scaled by a factor $1/n$ defines a lower bound for the asymptotic variance of nonparametric estimators of $\mathbb{E}_{P}(Y^{\boldsymbol{a}})$ for a given time-dependent adjustment set [9, 10]. We further define $\sigma^{2}_{\boldsymbol{a},\boldsymbol{Z}}(P)$ and $\sigma^{2}_{\Delta,\boldsymbol{Z}}(P)$ as the asymptotic variances of nonparametric efficient estimators of the counterfactual mean $\chi_{\boldsymbol{a}}(P;\mathcal{G})$ and of causal contrast $\Delta(P;\mathcal{G})$ under the time-dependent adjustment set $\boldsymbol{Z}$. Note that the latter can be obtained from the former using the functional delta method [11] or tricks delineated in [12]. Using this notation, the following two lemmas compare the asymptotic variance of nonparametric efficient estimators of causal contrasts between nested sufficient time-dependent adjustment sets.
Lemma 4.1.
(Inclusion of additional variables $(\boldsymbol{G},\boldsymbol{D})$ in a sufficient time-dependent adjustment set $\boldsymbol{Z}=(\boldsymbol{B},\boldsymbol{C})$) Consider a DAG $\mathcal{G}$ with vertex set $\boldsymbol{V}$. Let $\boldsymbol{A}\subset\boldsymbol{V}$ and $Y\in\boldsymbol{V}\setminus\boldsymbol{A}$. For $t=T,\dots,0$, define $\mathcal{G}_{t}$ as the subgraph of $\mathcal{G}$ constructed by applying Algorithm 1. Let $\boldsymbol{K}=(\boldsymbol{K}_{0},\boldsymbol{K}_{1},...,\boldsymbol{K}_{T})$ be a sufficient time-dependent adjustment set according to Definition 1 and let $\boldsymbol{B}_{t},\boldsymbol{G}_{t},\boldsymbol{C}_{t},\boldsymbol{D}_{t}\subseteq\{\bar{\boldsymbol{A}}_{t-1},\bar{\boldsymbol{K}}_{t}\}$, t=0,,Tt=0,...,T, be a partition of $\{\bar{\boldsymbol{A}}_{t-1},\bar{\boldsymbol{K}}_{t}\}$. Suppose $(\boldsymbol{B},\boldsymbol{C})=\{(\boldsymbol{B}_{0},\boldsymbol{C}_{0}),(\boldsymbol{B}_{1},\boldsymbol{C}_{1}),\dots,(\boldsymbol{B}_{T},\boldsymbol{C}_{T})\}$ is a sufficient time-dependent adjustment set according to Definition 2 that satisfies
$A_{t}\perp\!\!\!\perp_{\mathcal{G}_{t}}\boldsymbol{G}_{t},\boldsymbol{D}_{t}\ \mid\ \boldsymbol{B}_{t},\boldsymbol{C}_{t},\ for\ t=0,...,T.$ (4)
@@
Then, for any distribution $P$ compatible with the graph $\mathcal{G}$, $(\boldsymbol{G},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D})=\{(\boldsymbol{G}_{0},\boldsymbol{B}_{0},\boldsymbol{C}_{0},\boldsymbol{D}_{0}),(\boldsymbol{G}_{1},\boldsymbol{B}_{1},\boldsymbol{C}_{1},\boldsymbol{D}_{1}),\dots,(\boldsymbol{G}_{T},\boldsymbol{B}_{T},\boldsymbol{C}_{T},\boldsymbol{D}_{T})\}$ is also a sufficient time-dependent adjustment set according to Definition 2 and
  1. 1.
    $\sigma^{2}_{\boldsymbol{a},\boldsymbol{B},\boldsymbol{C}}(P)-\sigma^{2}_{\boldsymbol{a},\boldsymbol{G},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}}(P)\geq 0\ \ and\ \ \sigma^{2}_{\Delta,\boldsymbol{B},\boldsymbol{C}}(P)-\sigma^{2}_{\Delta,\boldsymbol{G},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}}(P)\geq 0.$
  2. 2.
    For any $\boldsymbol{G}^{\prime}\subseteq\boldsymbol{G}$ and $\boldsymbol{D}^{\prime}\subseteq\boldsymbol{D}$, ($\boldsymbol{G}^{\prime},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}^{\prime}$) is also sufficient time-dependent adjustment set according to Definition 2 and
    $\sigma^{2}_{\boldsymbol{a},\boldsymbol{B},\boldsymbol{C}}(P)\;\;\geq\;\;\sigma^{2}_{\boldsymbol{a},\boldsymbol{G}^{\prime},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}^{\prime}}(P)\;\;\geq\;\;\sigma^{2}_{\boldsymbol{a},\boldsymbol{G},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D}}(P).$
Lemma 11 of Rotnitzky and Smucler [5] and Lemma 4.1 rest on the same basic mechanism: supplementing a valid adjustment set with so-called precision variables that help predict the outcome but, conditional on the current adjustment set, are not associated with the treatment, cannot introduce bias and will not increase the asymptotic variance of the causal estimator.
Lemma 4.2.
(Exclusion of variables $(\boldsymbol{B},\boldsymbol{D})$ from a sufficient time-dependent adjustment set $\boldsymbol{Z}=(\boldsymbol{G},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D})$) Consider a DAG $\mathcal{G}$ with vertex set $\boldsymbol{V}$. Let $\boldsymbol{A}\subset\boldsymbol{V}$ and $Y\in\boldsymbol{V}\setminus\boldsymbol{A}$. For $t=T,\dots,0$, define $\mathcal{G}_{t}$ as the subgraph of $\mathcal{G}$ constructed by applying Algorithm 1. Let $\boldsymbol{K}=(\boldsymbol{K}_{0},\boldsymbol{K}_{1},...,\boldsymbol{K}_{T})$ be a sufficient time-dependent adjustment set according to Definition 1 and let $\boldsymbol{B}_{t},\boldsymbol{G}_{t},\boldsymbol{C}_{t},\boldsymbol{D}_{t}\subseteq\{\bar{\boldsymbol{A}}_{t-1},\bar{\boldsymbol{K}}_{t}\}$, t=0,,Tt=0,...,T, be a partition of $\{\bar{\boldsymbol{A}}_{t-1},\bar{\boldsymbol{K}}_{t}\}$. Suppose $(\boldsymbol{G},\boldsymbol{B},\boldsymbol{C},\boldsymbol{D})=\{(\boldsymbol{G}_{0},\boldsymbol{B}_{0},\boldsymbol{C}_{0},\boldsymbol{D}_{0}),(\boldsymbol{G}_{1},\boldsymbol{B}_{1},\boldsymbol{C}_{1},\boldsymbol{D}_{1}),\dots,(\boldsymbol{G}_{T},\boldsymbol{B}_{T},\boldsymbol{C}_{T},\boldsymbol{D}_{T})\}$ is a sufficient time-dependent adjustment set according to Definition 2 that satisfies
$Y\perp\!\!\!\perp_{\mathcal{G}_{T}}\boldsymbol{B}_{T},\boldsymbol{D}_{T}\ \mid\ \boldsymbol{G}_{T},\boldsymbol{C}_{T},{A}_{T}$ (5)